Assume the following information regarding U.S. and British currency values and 1-year rates: Spot Rate $1 = £0.8299 1-Year Forward Rate $1 = £0.8631 British 1-Year Interest Rate 7.0% U.S. 1-Year Interest Rate 5.0% Given this information, what is the yield (or profits) to a British investor who conducts covered interest arbitrage with 1,000,000 British pounds. Round to one decimal and give it as a percent. So, if you calculate the answer to be .0438, you would answer 4.4 for 4.4%. Leave off the percent sign. If you give the answer in profits, round to the nearest whole number. No commas or currency signs. Negative numbers should have a negative sign. Answer:
Answer #1

The steps to make a covered interest arbitrage profit are :

  • Borrow £1,000,000 at the British interst rate for 1 year. Convert into $ at the spot rate. $ received = £1,000,000 / 0.8299 = $1,204,964.45
  • Invest $1,204,964.45 at the US interest rate for 1 year. $ received after 1 year = $1,204,964.45 * (1 + 5%) = $1,265,212.68
  • Enter into a forward contract today to convert $1,265,212.68 into £ after 1 year. After 1 year, exercise the contract.  £ received = $1,265,212.68 * 0.8631 =  £1,092,005.06
  • Repay the original £ borrowed along with interest. £ to repay =  £1,000,0 * (1 + 7%) =  £1,070,000
  • Arbitrage profit = £ received after 1 year - £ to repay after 1 year
  • Arbitrage profit = £1,092,005.06 - £1,070,000
  • Arbitrage profit = £22,005

Yield to British investor = Arbitrage profit / £ borrowed

Yield to British investor = £22,005 / £1,000,000

Yield to British investor = 2.2%